OpenStudy (anonymous):

2x + 3y = 2 and -x + 2y =6

OpenStudy (anonymous):

\[2x+3y=2 ; 2x=2-3y; x= 1-\frac{ 3 }{ 2}y\]

OpenStudy (anonymous):

Your new equation is: \[-(1-\frac{ 3 }{ 2}y)+2y=6\]